ebhancoc:

Why is 2 + (−3) equal to −1?

katkit25:

you are technically saying 2 minus 1 so yeah

ZoeyBiocth16:

2+(-3)=-1 2-3=-1 -1=-1 this is because you have to keep the negative exponent which is with the biggest number.

dude:

Careful, its a co-efficient, not an exponent